OpenStudy (dtan5457):

Evaluate this equation if a=2

OpenStudy (alexandervonhumboldt2):

equation

OpenStudy (dtan5457):

\[a^2+3a^-4-a^0 \]

OpenStudy (dtan5457):

the 3a is ^-4

OpenStudy (alexandervonhumboldt2):

2*2+(3a)^-4-a^0

OpenStudy (alexandervonhumboldt2):

4-1/81(a^-4)-1=3-1/81a^-4

OpenStudy (dtan5457):

can you put that in equation mode i can't read it

OpenStudy (alexandervonhumboldt2):

\[3-\frac{ 1 }{ 81 }\left( a ^{-4} \right)\]

OpenStudy (dtan5457):

isn't 2^4=16

OpenStudy (dtan5457):

@DanJS

OpenStudy (dtan5457):

@jim_thompson5910

OpenStudy (alexandervonhumboldt2):

3^-4=1/81

OpenStudy (dtan5457):

if i plug in 2 3(2)^-4

OpenStudy (dtan5457):

what are my steps?

OpenStudy (alexandervonhumboldt2):

ahh watlet m redo

OpenStudy (dtan5457):

you mean 3^-4 times 2^-4?

OpenStudy (alexandervonhumboldt2):

2*2+(3*2)^-4

OpenStudy (dtan5457):

multiply 3 by 2 first?

OpenStudy (alexandervonhumboldt2):

if 3a is in -4 poweer then yes

OpenStudy (dtan5457):

oo it's the same result

OpenStudy (alexandervonhumboldt2):

4-1/1296

OpenStudy (dtan5457):

3 1/1296?

OpenStudy (alexandervonhumboldt2):

hmm no cant be

OpenStudy (alexandervonhumboldt2):

\[a ^{2}+(3a^{-4})\] is it correct?

OpenStudy (dtan5457):

there wasn't parenthesis but yes

OpenStudy (alexandervonhumboldt2):

k so (2^2)+(3*(2^-4))

OpenStudy (alexandervonhumboldt2):

4+3*1/16

OpenStudy (alexandervonhumboldt2):

67/16

OpenStudy (alexandervonhumboldt2):

4 3/16

OpenStudy (dtan5457):

you mean 3+3/16

OpenStudy (dtan5457):

-2^0=-1

OpenStudy (alexandervonhumboldt2):

4+3/16

OpenStudy (dtan5457):

it's 3

OpenStudy (dtan5457):

there is no 4

OpenStudy (alexandervonhumboldt2):

a=2 2^2=4

OpenStudy (dtan5457):

but let me ask you if it was (3 times 2)^-4 that would be 1/1256 right?

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (dtan5457):

alright good looks

OpenStudy (alexandervonhumboldt2):

o